OpenStudy (anonymous):

∫√(1+(e powered by 2)dx

OpenStudy (anonymous):

\[1+e^2\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

e to the x

OpenStudy (anonymous):

you're missing a square root

OpenStudy (anonymous):

cambell

OpenStudy (anonymous):

the contents are enclosed in a root sing

OpenStudy (campbell_st):

then its even easier as \[\sqrt{1 + e^2}\] is a constant \[\int\limits \sqrt{1 + e^2} dx = x \sqrt{1 + e^2} + c\]

OpenStudy (anonymous):

help me guys

OpenStudy (anonymous):

how do i then work it out

OpenStudy (anonymous):

^ what cambell said. you have a constant unless that last part is e^2x then you have something different

OpenStudy (anonymous):

its actually e raised to x

OpenStudy (anonymous):

not to 2

OpenStudy (campbell_st):

well \[\sqrt{1 + e^2} = 2.896\] its just a number.... you can plot it on the number line the the problem is just like \[\int\limits 5 dx = 5x + c\]

OpenStudy (anonymous):

\[\int\limits_{}^{}\sqrt{1+e ^{x}}\]

OpenStudy (campbell_st):

lol... now you have the question \[\int\limits \sqrt{1 + e^{2x}} dx\]

OpenStudy (anonymous):

only x not 2x

OpenStudy (anonymous):

dnt leave yet campbell i could really use your help with a couple qsns

OpenStudy (anonymous):

is it goin well

OpenStudy (campbell_st):

well then its integration by substitution let u = e^x du/dx = e^x du = e^xdx or dx = du/u \[\int\limits \sqrt{(1 + u)}/u du\] looks like a double substitution... good luck

OpenStudy (anonymous):

what then would be the final answer ?

OpenStudy (zarkon):

\[u=1+e^x\]